### v1

# Four special Poncelet triangle families about the incircle

Abstract.
We describe four special families of ellipse-inscribed Poncelet triangles about the incircle which maintain certain triangle centers stationary and which also display interesting conservations.
December 19, 2025
1. Introduction
In the spirit of the research done in [garcia2024-incircle], we study four special families of ellipse-inscribed Poncelet triangles about the incircle. Referring to Figure 1, these are:
  • focal-$X_{1}$: the caustic is centered on a focus of $\mathcal{E}$;
  • iso-$X_{2}$: caustic is centered on a first special point such that the barycenter $X_{2}$ is stationary;
  • focal-$X_{4}$: caustic is centered on a second special point such that the orthocenter $X_{4}$ is stationary on a focus of $\mathcal{E}$;
  • iso-$X_{7}$: caustic is centered on a third special point such that the Gergonne point $X_{7}$ is stationary.
In the calculations below, $R$, $r$, $l_{i},i=1,2,3$ refer to a triangle’s circumradius, inradius, and sidelengths, respectively.
Figure 1. Four families of ellipse-inscribed Poncelet triangles about the incircle. top-left: focal-$X_{1}$ ($C$ at a focus of $\mathcal{E}$); top-right: iso-$X_{2}$ ($C$ on minor axis of $\mathcal{E}$, barycenter $X_{2}$ stationary); bot.-left: focal-$X_{4}$ ($X_{4}$ is a focus of $\mathcal{E}$); bot.-right: iso-$X_{7}$ ($C$ on major axis of $\mathcal{E}$, $X_{7}$ stationary). Video: youtu.be/nsHDfX6_7mA
2. Focal-$X_{1}$
Referring to Figure 1 (top left):
Proposition 1.
The caustic centered at $C_{1}=[c,0]$ admits a Poncelet family of triangles with $r_{1}$ given by:
$r_{1}={\frac{{b}^{2}}{{c}^{2}}\left(\sqrt{{a}^{2}+{c}^{2}}-a\right)}\cdot$
Proof.
This follows from [garcia2024-incircle, Prop. 2]: the radius $r$ for a circular caustic (of Poncelet triangles) with center at $x_{c},y_{c}]$ is given by:
$r=\frac{b\sqrt{a^{4}-c^{2}x_{c}^{2}}-a\sqrt{b^{4}+c^{2}y_{c}^{2}}}{c^{2}}\cdot$
Observation 1.
The focal-$X_{1}$ family is the polar image of the vertices of Chapple’s Porism (a porism of triangles interscribed between two circles [odehnal2011-poristic]), with respect to the circumcircle, i.e., it is its tangential triangle. The former’s incenter $X_{1}$ coincides with the latter’s circumcenter $X_{3}$.
Bicentric $n$-gons (a generalization of Chapple’s porism) conserve the sum of cosines of their internal angles $\theta_{i}$ [roitman2021-bicentric]. On the other hand, the polar family with respect to the circumcircle conserves $\sum_{i=1}^{n}{\sin(\theta_{i}/2)}$, for all $n\geq 3$ [bellio2022-parabola-inscribed, prop.26].
Proposition 2.
For the focal-$X_{1}$ family, the invariant sum of half-angle sines is given by:
$\sum_{i=1}^{3}\sin\frac{\theta_{i}}{2}=\frac{c^{2}-a^{2}+a\sqrt{a^{2}+c^{2}}}{c^{2}}\cdot$
3. Iso-$X_{2}$
Referring to Figure 1 (top right), it follows from the expressions for the center $C_{2}$ and semi-axis lengths $a_{2},b_{2}$ for the elliptic locus of the barycenter $X_{2}$ [garcia2024-incircle, Prop. 3] that:
Proposition 3.
For a circular caustic with center $C_{2}=\left[0,\frac{c\,b}{2a}\right]$ and radius $r_{2}=\frac{b}{2}$, the barycenter $X_{2}$ is stationary at $\left[0,\frac{c\,b}{3a}\right]$.
Corollary 1.
Over iso-$X_{2}$ triangles, Nagel’s point $X_{8}$ is stationary at $E$’s center.
Proof.
Direct from the location of the stationary barycenter and the fact that $|X_{1}X_{8}|=3|X_{1}X_{2}|$ [mw, Incenter, eqn.11]. In fact, $|X_{1}X_{8}|=3(c\,b/(2a)-c\,b/(3a))=c\,b/(2a)=|C_{2}|$. ∎
Corollary 2.
The Spieker center $X_{10}$, incenter of the medial triangle and the midpoint of the incenter $X_{1}$ and Nagel’s point $X_{8}$ [etc] will be stationary on $E$’s minor axis.
Observation 2.
Since the barycenter $X_{2}$ is stationary, the family conserves $|X_{1}X_{2}|$, which for any triangle with sidelengths $l_{i}$ is given by [mw, Triangle Centroid, eqn.11]:
$|X_{1}-X_{2}|^{2}=-\frac{\sum{l_{i}^{3}}+9\prod{l_{i}}-2\left(l_{2}l_{1}^{2}+l_{3}l_{1}^{2}+l_{2}^{2}l_{1}+l_{3}^{2}l_{1}+l_{2}l_{3}^{2}+l_{2}^{2}l_{3}\right)}{9\sum{l_{i}}}\raisebox{1.29167pt}{.}$
Definition 1.
The MacBeath Poncelet family, shown in Figure 5 (left), are triangles inscribed in a circle, and circumscribing the MacBeath inconic, whose foci are the circumcenter $X_{3}$ and the orthocenter $X_{4}$, Its center is the Euler center $X_{5}$, the midpoint of the circumcenter $X_{3}$ and the orthocenter $X_{4}$ [mw, MacBeath inconic].
Lemma 1.
The barycenter $X_{2}$ of the MacBeath family is stationary and lies on the major axis of the inconic.
Proof.
The foci of the caustic are the circumcenter $X_{3}$ and the orthocenter $X_{4}$, therefore the barycenter $X_{2}$ must be fixed, $|X_{3}X_{2}|=|X_{3}X_{4}|/3$ [mw, Triangle Centroid, eqn.(13)]. Consider also that MacBeath family are the intouch triangles of bicentric polygons (Chapple’s porism). The the barycenter $X_{2}$ of the intouch triangle is the Weill point $X_{354}$ of the reference [etc], shown to be stationary over the bicentric family [odehnal2011-poristic, Thm. 4.1]. Since the barycenter $X_{2}$ lies on the Euler line, which also contains $X_{3},X_{4}$, it must lie on the inconic’s axis. ∎
Definition 2 (Isogonal conjugate).
The isogonal conjugate $P^{\dagger}$ of a point $P$ with respect to a triangle $T$ is the point of concurrence of the cevians of $P$ reflected upon the angle bisectors. If the barycentrics of $P$ are $z_{1}:z_{2}:z_{3}]$, $P^{\dagger}=[l_{1}^{2}/z_{1}:l_{2}^{2}/z_{2}:l_{3}^{2}/z_{3}]$, where the $l_{i}$ are the sidelengths.
Proposition 4.
Given an outer ellipse $\mathcal{E}$ with axes $(a,b)$ centered on $O$ and a point $O_{c}$ in the interior of a concentric, axis-aligned ellipse with axes $(a/2,b/2)$, there is a caustic $\mathcal{K}$ which admits a Poncelet triangle family. Over this family, the barycenter $X_{2}$ is stationary, with the latter lying on $OO_{c}$.
Proof.
Referring to Figure 2, let $\mathcal{A}$ be an affine transform that sends an outer Poncelet conic $\mathcal{E}$ centered at $O$ (left) to a circle $\mathcal{E}^{\prime}$ (right), centered at $O^{\prime}=\mathcal{A}(O)$ (conic centers are equivariant under affinities). Since $O_{c}$ is as specified, $O_{c}^{\prime}=\mathcal{A}(O_{c})$ will lie within a circle concentric with $\mathcal{E}^{\prime}$ and of half its radius.
Let $\mathcal{K}^{\prime}$ be the unique Poncelet caustic with a first focus at $O^{\prime}$ (the circumcenter $X_{3}$ of its triangles) and the other at the reflection of $\mathcal{O}^{\prime}$ about $O_{c}^{\prime}=\mathcal{A}(O_{c})$. Since foci of a triangle’s inconic are an isogonal conjugate pair [beluhov2007-isogonal-foci] (see also Definition 2), the second focus must be the fixed orthocenter $X_{4}$ of the family (also within $\mathcal{E}^{\prime}$), i.e., this is the MacBeath inconic.
Let $\mathcal{K}$ be the image of $\mathcal{K}^{\prime}$ under $\mathcal{A}^{-1}$. The the barycenter $X_{2}$ of the family in $(\mathcal{E},\mathcal{K})$ will be stationary because (i) by Lemma 1, the MacBeath’s centroid is stationary, and (ii) the centroid is the only triangle center which is equivariant with respect to affinities [etc]. The barycenter $X_{2}$ must lie on $OC$ because affinities preserve collinearities. Notice it will nevertheless not lie on the major axis of $\mathcal{K}$ (foci are not equivariant). ∎
Figure 2. Constructions for the proof of Proposition 4: an affine transformation exists such that the outer ellipse (right) is sent to a circle (right). The dashed half-size ellipse in the left (half radius circle on the right) delimits the region for acceptable $O_{c}$ ($O_{c}^{\prime}$ on the right). Video: https://youtu.be/yDz_XopFw2Y
Referring to Figure 3, experimental evidence suggests that for MacBeath-like Poncelet families of $n$-gons, $n>3$, i.e., they are circle-inscribed and the caustic has a focus at the center of said circle, both vertex and area centroids $C_{0}$ and $C_{2}$ remain stationary on the caustic axis. When $n=4$, $C_{0}$ lies at the caustic’s center.
Corollary 3.
Given an ellipse $E$ centered at $O$ and a point $C$ in its interior, there is a caustic centered on $C$ for which both $c_{0}$ and $c_{2}$ are stationary on $OC$.
Indeed, over such families, the perimeter centroid, in general not expected to sweep a conic over Poncelet [sergei2016-com], is experimentally found to sweep an ellipse with major axis identical to the caustic’s.
Figure 3. MacBeath-like Poncelet families of quadrilaterals (left) and pentagons (right). In both cases The vertex and area centroids $C_{0},C_{2}$ are stationary on the caustic’s major axis, whereas the perimeter centroid $C_{1}$ sweeps an ellipse. In the former case, $C_{0}$ is located at the center of the caustic. Video: youtu.be/e-iOAAFU1H4
4. Focal-$X_{4}$
Referring to Figure 1 (bottom left):
Proposition 5.
The orthocenter $X_{4}$ is stationary at a focus $f=[\pm c,0]$ of $E$ for the circular caustic with center and radius given by:
$C_{4}=\left[\pm\frac{a^{2}c}{2a^{2}-c^{2}},0\right],\,\,\,r_{4}=\frac{a(a^{2}-c^{2})}{2a^{2}-c^{2}}\raisebox{1.29167pt}{.}$
Referring to Figure 4:
Observation 3.
For the focal-$X_{4}$ family, the center (resp. other focus) of $\mathcal{E}$ is $X_{7952}$ (resp. $X_{18283}$). $\mathcal{L}_{3}$ is an ellipse with a focus at $E$’s center. $\mathcal{L}_{20}$ is twice as large, with a focus on $\mathcal{E}$’s distal focus.
Figure 4. The focal-$X_{4}$ family: the outer ellipse $E$ has foci on the orthocenter $X_{4}$ and $X_{18283}$ and center on $X_{7952}$. Also shown are the loci of the circumcenter $X_{3}$, with a focus on the latter center, and that of the de Longchamps point $X_{20}$ (a reflection of the orthocenter $X_{4}$ on the circumcenter $X_{3}$), so twice as big, and with a focus on the other focus of $E$. Video: youtu.be/x9ce56vvDNI
Definition 3.
A triangle’s polar circle is centered on the orthocenter $X_{4}$ and has squared radius given by [mw, Polar circle]:
$r_{pol}^{2}=4R^{2}-\frac{\sum{l_{i}^{2}}}{2}\raisebox{1.29167pt}{.}$
This quantity is positive (resp. negative) for obtuse (resp. acute triangles).
Proposition 6.
The focal-$X_{4}$ family conserves the (negative) squared radius of its polar circle at $r^{2}_{pol}=-b^{4}/(a^{2}+b^{2}).$
Proof.
The incenter-orthocenter squared distance is given by [mw, Incenter, eqn.6]:
$|X_{1}-X_{4}|^{2}=2r^{2}+4R^{2}-\frac{\sum{l_{i}^{2}}}{2}=2r^{2}+r_{pol}^{2}\,\cdot$
Since for this family $|X_{1}X_{4}|$ is fixed as is the inradius $r$. The expression was obtained by CAS-based simplification. ∎
Let $a,b$ denote the semiaxis’ lengths of the MacBeath inconic.
Lemma 2.
Over the MacBeath family, the sum of squared sidelengths, the sum of double-angle cosines, and the product of cosines are conserved. There are given by:
$\sum{l_{i}^{2}}$ $=9R^{2}-|X_{3}-X_{4}|^{2}=32a^{2}+4b^{2},$
$\sum{\cos(2\theta_{i})}$ $=\frac{c^{2}-3a^{2}}{2a^{2}}\raisebox{1.29167pt}{,}$
$\prod{\cos\theta_{i}}$ $=\frac{\sum{l_{i}^{2}}}{8R^{2}}-1=\frac{b^{2}}{8a^{2}}\cdot$
Proof.
The first relation is direct from [mw, Circumcenter, eqn.6], noting that (i) the circumradius $R$ is fixed and (ii) the caustic foci $X_{3},X_{4}$ are stationary. The second one is obtained via CAS-simplification. The third one is direct from [mw, Circumradius, eqn.5]. ∎
Referring to Figure 5 (right):
Definition 4.
The dual Poncelet family is interscribed between two dual ellipses [walker1950-curves]. They are also homothetic if one is rotated by 90-degrees.
Figure 5. left: the ‘MacBeath’ family. Live: bit.ly/4eVqLAS; right: the ‘dual’ family. Live: bit.ly/4eSQSsd
Proposition 7.
The (i) MacBeath and (ii) dual families conserve $r_{pol}^{2}<0$. These are given by:
$r_{pol,macbeath}^{2}$ $=\frac{|X_{3}-X_{4}|^{2}-R^{2}}{2}=-2b^{2}\raisebox{1.29167pt}{,}$
$r_{pol,dual}^{2}$ $=-\frac{a^{2}b^{2}}{a^{2}+b^{2}}\,\cdot$
Proof.
The first is direct from $|X_{3}-X_{4}|^{2}=9R^{2}-\sum{l_{i}^{2}}=2r_{pol}^{2}+R^{2}$ [mw, Orthocenter, eqn.14], noting that the MacBeath circumradius $R$ is fixed and the caustic foci on the circumcenter $X_{3}$ and orthocenter $X_{4}$ are stationary. For the dual family, the orthocenter $X_{4}$ is stationary at the common center. Given the first relation it would be sufficient to prove that $|X_{3}(t)|^{2}-R^{2}(t)$ is constant. The final expression is obtained via CAS simplification. ∎
Observation 4.
For the dual family, $\mathcal{L}_{3}$ is an ellipse homothetic to $\mathcal{E}$ with factor $c^{2}/(2(a^{2}+b^{2}))$.
Conjecture 1.
A Poncelet triangle family maintains the orthocenter $X_{4}$ stationary only if the conics are configured as focal-$X_{4}$, MacBeath or Dual.
5. Iso-$X_{7}$
Referring to Figure 1 (bottom right):
Proposition 8.
For the circular caustic with center $C_{7}=\left[k_{7}/(2a),0\right]$ and radius $r_{7}=b^{2}/(2a)$, the Gergonne point will be stationary at:
$X_{7}=\left[\frac{2ak_{7}}{4a^{2}-b^{2}},0\right],$
where $k_{7}=\sqrt{4a^{4}-5a^{2}b^{2}+b^{4}}$.
Observation 5.
The iso-$X_{7}$ family is the polar image of the Brocard porism [reznik2022-brocard-converging, shail1996-brocard] with respect to the circumcircle, i.e., its tangential triangle. The former’s stationary Gergonne point $X_{7}$ coincides with the stationary symmedian point $X_{6}$ of the Brocard porism [garcia2020-family-ties].
Proposition 9.
The family conserves the sum of half-angle tangents, and this is given by:
$\sum_{i=1}^{3}{\tan\frac{\theta_{i}}{2}}=\frac{\sqrt{4\,{a}^{2}-{b}^{2}}}{a}\raisebox{1.29167pt}{.}$
The $N>3$ generalization of the Brocard porism is known as the Harmonic family, studied in [roitman2022-harmonic].
Conjecture 2.
The sum of half-angle tangents is also conserved for the polar image of the Harmonic family, for all $N>3$.
Since the Gergonne center $X_{7}$ is stationary, the quantity $|X_{1}X_{7}|$ will also be conserved. For a generic triangle with semiperimeter $s$, this is given by [queiroz2012-gergonne, Corollary 4.2]:
$|X_{1}-X_{7}|^{2}=r^{2}\left(1-\frac{3s^{2}}{(r+4R)^{2}}\right)\raisebox{1.29167pt}{.}$
Proposition 10.
Over Iso-$X_{7}$ triangles, the previous expression reduces to: $|X_{1}-X_{7}|^{2}=b^{4}c^{2}/\left(4a^{2}(4a^{2}-b^{2})\right)$.
Definition 5.
The Adams circle of a triangle is centered on the incenter $X_{1}$ and has radius $R_{A}$ given by [mw, Adams’ Circle]: $R_{A}=\frac{r\sqrt{\rho^{2}-l_{1}l_{2}l_{3}s-\rho s^{2}}}{\rho-s^{2}}$, where $\rho=l_{1}l_{2}+l_{2}l_{3}+l_{3}l_{1}$ and $s$ is the semi-perimeter.
Proposition 11.
The Iso-$X_{7}$ family conserves $R_{A}$ and its value is given by:
$R_{A}=\frac{b^{2}}{2a}\sqrt{\frac{5a^{2}-b^{2}}{4a^{2}-b^{2}}}\,\raisebox{1.29167pt}{.}$