### v2

# Four special Poncelet triangle families about the incircle

Abstract.
We describe four special families of ellipse-inscribed Poncelet triangles about the incircle which maintain certain triangle centers stationary and which also display interesting conservations.
December 22, 2025
1. Introduction
As special cases to [garcia2024-incircle], we introduce four special families of Poncelet triangles inscribed in an ellipse $\mathcal{E}$ and circumscribing their fixed incircle, let $C$ be its center, i.e., the incenter $X_{1}$ (triangle center notation $X_{k}$ are after Kimberling [etc]). Referring to Figure 1, the four Poncelet families are:
  • focal-$X_{1}$: the caustic is centered on a focus of $\mathcal{E}$;
  • iso-$X_{2}$: caustic is centered on a special point on $\mathcal{E}$’s minor axis such that the barycenter $X_{2}$ is stationary;
  • focal-$X_{4}$: caustic is centered on a special point $\mathcal{E}$’s major axis such that the orthocenter $X_{4}$ is stationary on a focus of $\mathcal{E}$;
  • iso-$X_{7}$: caustic is centered on a another special point on $\mathcal{E}$’s major axis such that the Gergonne point $X_{7}$ is stationary.
In the sections below, $R$, $r$, $l_{i},i=1,2,3$ refer to a triangle’s circumradius, inradius, and sidelengths, respectively, see [mw] for definitions.
Figure 1. Four families of ellipse-inscribed Poncelet triangles about the incircle. top-left: focal-$X_{1}$ ($C$ at a focus of $\mathcal{E}$); top-right: iso-$X_{2}$ ($C$ on minor axis of $\mathcal{E}$, barycenter $X_{2}$ stationary); bot.-left: focal-$X_{4}$ ($X_{4}$ is a focus of $\mathcal{E}$); bot.-right: iso-$X_{7}$ ($C$ on major axis of $\mathcal{E}$, $X_{7}$ stationary). Video: youtu.be/nsHDfX6_7mA
2. Focal-$X_{1}$
Let $\mathcal{E}$ be given by $(x/a)^{2}+(y/b)^{2}=0$. Referring to Figure 1 (top left):
Proposition 1.
The caustic centered at $C_{1}=[c,0]$ admits a Poncelet family of triangles with $r_{1}$ given by:
$r_{1}={\frac{{b}^{2}}{{c}^{2}}\left(\sqrt{{a}^{2}+{c}^{2}}-a\right)}\cdot$
Proof.
This follows from [garcia2024-incircle, Prop. 2]: the radius $r$ for a circular caustic (of Poncelet triangles) with center at $x_{c},y_{c}]$ is given by:
$r=\frac{b\sqrt{a^{4}-c^{2}x_{c}^{2}}-a\sqrt{b^{4}+c^{2}y_{c}^{2}}}{c^{2}}\cdot\qed$
Observation 1.
The focal-$X_{1}$ family is the polar image of the vertices of Chapple’s Porism (a porism of triangles interscribed between two circles [odehnal2011-poristic]), with respect to the circumcircle, i.e., it is its tangential triangle. The former’s incenter $X_{1}$ coincides with the latter’s circumcenter $X_{3}$.
Bicentric $n$-gons (a generalization of Chapple’s porism) conserve the sum of cosines of their internal angles $\theta_{i}$ [roitman2021-bicentric]. On the other hand, the polar family with respect to the circumcircle conserves $\sum_{i=1}^{n}{\sin(\theta_{i}/2)}$, for all $n\geq 3$ [bellio2022-parabola-inscribed, prop.26].
Proposition 2.
For the focal-$X_{1}$ family, the sum of half-angle sines is invariant and given by:
$\sum_{i=1}^{3}\sin\frac{\theta_{i}}{2}=\frac{c^{2}-a^{2}+a\sqrt{a^{2}+c^{2}}}{c^{2}}\cdot$
3. Iso-$X_{2}$
Referring to Figure 1 (top right), it follows from the expressions for the center $C_{2}$ and semi-axis lengths $a_{2},b_{2}$ for the elliptic locus of the barycenter $X_{2}$ [garcia2024-incircle, Prop. 3] that:
Proposition 3.
For a circular caustic with center $C_{2}=\left[0,\frac{c\,b}{2a}\right]$ and radius $r_{2}=\frac{b}{2}$, the barycenter $X_{2}$ is stationary at $\left[0,\frac{c\,b}{3a}\right]$.
Corollary 1.
Over iso-$X_{2}$ triangles, Nagel’s point $X_{8}$ is stationary at $E$’s center.
Proof.
Direct from the location of the stationary barycenter and the fact that $|X_{1}X_{8}|=3|X_{1}X_{2}|$ [mw, Incenter, eqn.11]. In fact, $|X_{1}X_{8}|=3(c\,b/(2a)-c\,b/(3a))=c\,b/(2a)=|C_{2}|$. ∎
Corollary 2.
The Spieker center $X_{10}$, incenter of the medial triangle and the midpoint of the incenter $X_{1}$ and Nagel’s point $X_{8}$ [etc] will be stationary on $E$’s minor axis.
Observation 2.
Since the barycenter $X_{2}$ is stationary, the family conserves $|X_{1}X_{2}|$, which for any triangle with sidelengths $l_{i}$ is given by [mw, Triangle Centroid, eqn.11]:
$|X_{1}-X_{2}|^{2}=-\frac{\sum{l_{i}^{3}}+9\prod{l_{i}}-2\left(l_{2}l_{1}^{2}+l_{3}l_{1}^{2}+l_{2}^{2}l_{1}+l_{3}^{2}l_{1}+l_{2}l_{3}^{2}+l_{2}^{2}l_{3}\right)}{9\sum{l_{i}}}\raisebox{1.29167pt}{.}$
Definition 1.
The MacBeath Poncelet family, shown in Figure 5 (left), are triangles inscribed in a circle, and circumscribing the MacBeath inconic, whose foci are the circumcenter $X_{3}$ and the orthocenter $X_{4}$, Its center is the Euler center $X_{5}$, the midpoint of the circumcenter $X_{3}$ and the orthocenter $X_{4}$ [mw, MacBeath inconic].
Lemma 1.
The barycenter $X_{2}$ of the MacBeath family is stationary and lies on the major axis of the inconic.
Proof.
The foci of the caustic are the circumcenter $X_{3}$ and the orthocenter $X_{4}$, therefore the barycenter $X_{2}$ must be fixed, $|X_{3}X_{2}|=|X_{3}X_{4}|/3$ [mw, Triangle Centroid, eqn.(13)]. Consider also that MacBeath family are the intouch triangles of bicentric polygons (Chapple’s porism). The the barycenter $X_{2}$ of the intouch triangle is the Weill point $X_{354}$ of the reference [etc], shown to be stationary over the bicentric family [odehnal2011-poristic, Thm. 4.1]. Since the barycenter $X_{2}$ lies on the Euler line, which also contains $X_{3},X_{4}$, it must lie on the inconic’s axis. ∎
Definition 2 (Isogonal conjugate).
The isogonal conjugate $P^{\dagger}$ of a point $P$ with respect to a triangle $T$ is the point of concurrence of the cevians of $P$ reflected upon the angle bisectors. If the barycentrics of $P$ are $z_{1}:z_{2}:z_{3}]$, $P^{\dagger}=[l_{1}^{2}/z_{1}:l_{2}^{2}/z_{2}:l_{3}^{2}/z_{3}]$, where the $l_{i}$ are the sidelengths.
Proposition 4.
Given an outer ellipse $\mathcal{E}$ with axes $(a,b)$ centered on $O$ and a point $O_{c}$ in the interior of a concentric, axis-aligned ellipse with axes $(a/2,b/2)$, there is a caustic $\mathcal{K}$ which admits a Poncelet triangle family. Over this family, the barycenter $X_{2}$ is stationary, with the latter lying on $OO_{c}$.
Proof.
Referring to Figure 2, let $\mathcal{A}$ be an affine transform that sends an outer Poncelet conic $\mathcal{E}$ centered at $O$ (left) to a circle $\mathcal{E}^{\prime}$ (right), centered at $O^{\prime}=\mathcal{A}(O)$ (conic centers are equivariant under affinities). Since $O_{c}$ is as specified, $O_{c}^{\prime}=\mathcal{A}(O_{c})$ will lie within a circle concentric with $\mathcal{E}^{\prime}$ and of half its radius.
Let $\mathcal{K}^{\prime}$ be the unique Poncelet caustic with a first focus at $O^{\prime}$ (the circumcenter $X_{3}$ of its triangles) and the other at the reflection of $\mathcal{O}^{\prime}$ about $O_{c}^{\prime}=\mathcal{A}(O_{c})$. Since foci of a triangle’s inconic are an isogonal conjugate pair [beluhov2007-isogonal-foci] (see also Definition 2), the second focus must be the fixed orthocenter $X_{4}$ of the family (also within $\mathcal{E}^{\prime}$), i.e., this is the MacBeath inconic.
Let $\mathcal{K}$ be the image of $\mathcal{K}^{\prime}$ under $\mathcal{A}^{-1}$. The the barycenter $X_{2}$ of the family in $(\mathcal{E},\mathcal{K})$ will be stationary because (i) by Lemma 1, the MacBeath’s centroid is stationary, and (ii) the centroid is the only triangle center which is equivariant with respect to affinities [etc]. The barycenter $X_{2}$ must lie on $OC$ because affinities preserve collinearities. Notice it will nevertheless not lie on the major axis of $\mathcal{K}$ (foci are not equivariant). ∎
Figure 2. Constructions for the proof of Proposition 4: an affine transformation exists such that the outer ellipse (right) is sent to a circle (right). The dashed half-size ellipse in the left (half radius circle on the right) delimits the region for acceptable $O_{c}$ ($O_{c}^{\prime}$ on the right). Video: https://youtu.be/yDz_XopFw2Y
Referring to Figure 3, experimental evidence suggests that for MacBeath-like Poncelet families of $n$-gons, $n>3$, i.e., they are circle-inscribed and the caustic has a focus at the center of said circle, both vertex and area centroids $C_{0}$ and $C_{2}$ remain stationary on the caustic axis. When $n=4$, $C_{0}$ lies at the caustic’s center.
Corollary 3.
Given an ellipse $E$ centered at $O$ and a point $C$ in its interior, there is a caustic centered on $C$ for which both $c_{0}$ and $c_{2}$ are stationary on $OC$.
Indeed, over such families, the perimeter centroid, in general not expected to sweep a conic over Poncelet [sergei2016-com], is experimentally found to sweep an ellipse with major axis identical to the caustic’s.
Figure 3. MacBeath-like Poncelet families of quadrilaterals (left) and pentagons (right). In both cases The vertex and area centroids $C_{0},C_{2}$ are stationary on the caustic’s major axis, whereas the perimeter centroid $C_{1}$ sweeps an ellipse. In the former case, $C_{0}$ is located at the center of the caustic. Video: youtu.be/e-iOAAFU1H4
4. Focal-$X_{4}$
Referring to Figure 1 (bottom left):
Proposition 5.
The orthocenter $X_{4}$ is stationary at a focus $f=[\pm c,0]$ of $E$ for the circular caustic with center and radius given by:
$C_{4}=\left[\pm\frac{a^{2}c}{2a^{2}-c^{2}},0\right],\,\,\,r_{4}=\frac{a(a^{2}-c^{2})}{2a^{2}-c^{2}}\raisebox{1.29167pt}{.}$
Referring to Figure 4:
Observation 3.
For the focal-$X_{4}$ family, the center (resp. other focus) of $\mathcal{E}$ is $X_{7952}$ (resp. $X_{18283}$). $\mathcal{L}_{3}$ is an ellipse with a focus at $E$’s center. $\mathcal{L}_{20}$ is twice as large, with a focus on $\mathcal{E}$’s distal focus.
Figure 4. The focal-$X_{4}$ family: the outer ellipse $E$ has foci on the orthocenter $X_{4}$ and $X_{18283}$ and center on $X_{7952}$. Also shown are the loci of the circumcenter $X_{3}$, with a focus on the latter center, and that of the de Longchamps point $X_{20}$ (a reflection of the orthocenter $X_{4}$ on the circumcenter $X_{3}$), so twice as big, and with a focus on the other focus of $E$. Video: youtu.be/x9ce56vvDNI
Definition 3.
A triangle’s polar circle is centered on the orthocenter $X_{4}$ and has squared radius given by [mw, Polar circle]:
$r_{pol}^{2}=4R^{2}-\frac{\sum{l_{i}^{2}}}{2}\raisebox{1.29167pt}{.}$
This quantity is positive (resp. negative) for obtuse (resp. acute triangles).
Proposition 6.
The focal-$X_{4}$ family conserves the (negative) squared radius of its polar circle at $r^{2}_{pol}=-b^{4}/(a^{2}+b^{2}).$
Proof.
The incenter-orthocenter squared distance is given by [mw, Incenter, eqn.6]:
$|X_{1}-X_{4}|^{2}=2r^{2}+4R^{2}-\frac{\sum{l_{i}^{2}}}{2}=2r^{2}+r_{pol}^{2}\,\cdot$
Since for this family $|X_{1}X_{4}|$ is fixed as is the inradius $r$. The expression was obtained by CAS-based simplification. ∎
Let $a,b$ denote the semiaxis’ lengths of the MacBeath inconic.
Lemma 2.
Over the MacBeath family, the sum of squared sidelengths, the sum of double-angle cosines, and the product of cosines are conserved. There are given by:
$\sum{l_{i}^{2}}$ $=9R^{2}-|X_{3}-X_{4}|^{2}=32a^{2}+4b^{2},$
$\sum{\cos(2\theta_{i})}$ $=\frac{c^{2}-3a^{2}}{2a^{2}}\raisebox{1.29167pt}{,}$
$\prod{\cos\theta_{i}}$ $=\frac{\sum{l_{i}^{2}}}{8R^{2}}-1=\frac{b^{2}}{8a^{2}}\cdot$
Proof.
The first relation is direct from [mw, Circumcenter, eqn.6], noting that (i) the circumradius $R$ is fixed and (ii) the caustic foci $X_{3},X_{4}$ are stationary. The second one is obtained via CAS-simplification. The third one is direct from [mw, Circumradius, eqn.5]. ∎
Referring to Figure 5 (right):
Definition 4.
The dual Poncelet family is interscribed between two dual ellipses [walker1950-curves]. They are also homothetic if one is rotated by 90-degrees.
Figure 5. left: the ‘MacBeath’ family. Live: bit.ly/4eVqLAS; right: the ‘dual’ family. Live: bit.ly/4eSQSsd
Proposition 7.
The (i) MacBeath and (ii) dual families conserve $r_{pol}^{2}<0$. These are given by:
$r_{pol,macbeath}^{2}$ $=\frac{|X_{3}-X_{4}|^{2}-R^{2}}{2}=-2b^{2}\raisebox{1.29167pt}{,}$
$r_{pol,dual}^{2}$ $=-\frac{a^{2}b^{2}}{a^{2}+b^{2}}\,\cdot$
Proof.
The first is direct from $|X_{3}-X_{4}|^{2}=9R^{2}-\sum{l_{i}^{2}}=2r_{pol}^{2}+R^{2}$ [mw, Orthocenter, eqn.14], noting that the MacBeath circumradius $R$ is fixed and the caustic foci on the circumcenter $X_{3}$ and orthocenter $X_{4}$ are stationary. For the dual family, the orthocenter $X_{4}$ is stationary at the common center. Given the first relation it would be sufficient to prove that $|X_{3}(t)|^{2}-R^{2}(t)$ is constant. The final expression is obtained via CAS simplification. ∎
Observation 4.
For the dual family, $\mathcal{L}_{3}$ is an ellipse homothetic to $\mathcal{E}$ with factor $c^{2}/(2(a^{2}+b^{2}))$.
Conjecture 1.
A Poncelet triangle family maintains the orthocenter $X_{4}$ stationary only if the conics are configured as focal-$X_{4}$, MacBeath or Dual.
5. Iso-$X_{7}$
Referring to Figure 1 (bottom right):
Proposition 8.
For the circular caustic with center $C_{7}=\left[k_{7}/(2a),0\right]$ and radius $r_{7}=b^{2}/(2a)$, the Gergonne point will be stationary at:
$X_{7}=\left[\frac{2ak_{7}}{4a^{2}-b^{2}},0\right],$
where $k_{7}=\sqrt{4a^{4}-5a^{2}b^{2}+b^{4}}$.
Observation 5.
The iso-$X_{7}$ family is the polar image of the Brocard porism [reznik2022-brocard-converging, shail1996-brocard] with respect to the circumcircle, i.e., its tangential triangle. The former’s stationary Gergonne point $X_{7}$ coincides with the stationary symmedian point $X_{6}$ of the Brocard porism [garcia2020-family-ties].
Proposition 9.
The family conserves the sum of half-angle tangents, and this is given by:
$\sum_{i=1}^{3}{\tan\frac{\theta_{i}}{2}}=\frac{\sqrt{4\,{a}^{2}-{b}^{2}}}{a}\raisebox{1.29167pt}{.}$
The $N>3$ generalization of the Brocard porism is known as the Harmonic family, studied in [roitman2022-harmonic].
Conjecture 2.
The sum of half-angle tangents is also conserved for the polar image of the Harmonic family, for all $N>3$.
Since the Gergonne center $X_{7}$ is stationary, the quantity $|X_{1}X_{7}|$ will also be conserved. For a generic triangle with semiperimeter $s$, this is given by [queiroz2012-gergonne, Corollary 4.2]:
$|X_{1}-X_{7}|^{2}=r^{2}\left[1-\frac{3s^{2}}{(r+4R)^{2}}\right]\raisebox{1.29167pt}{.}$
Proposition 10.
Over Iso-$X_{7}$ triangles, the previous expression reduces to: $|X_{1}-X_{7}|^{2}=b^{4}c^{2}/\left(4a^{2}(4a^{2}-b^{2})\right)$.
Definition 5.
The Adams circle of a triangle is centered on the incenter $X_{1}$ and has radius $R_{A}$ given by [mw, Adams’ Circle]: $R_{A}=\frac{r\sqrt{\rho^{2}-l_{1}l_{2}l_{3}s-\rho s^{2}}}{\rho-s^{2}}$, where $\rho=l_{1}l_{2}+l_{2}l_{3}+l_{3}l_{1}$ and $s$ is the semi-perimeter.
Proposition 11.
The Iso-$X_{7}$ family conserves $R_{A}$ and its value is given by:
$R_{A}=\frac{b^{2}}{2a}\sqrt{\frac{5a^{2}-b^{2}}{4a^{2}-b^{2}}}\,\raisebox{1.29167pt}{.}$